OpenStudy (anonymous):

Find P(rolling 2 or 5) with one number cube. A. 1/2 B. 1/3 C. 1/6 D. 1

jimthompson5910 (jim_thompson5910):

How many sides are on a number cube?

jimthompson5910 (jim_thompson5910):

Think of a die

OpenStudy (anonymous):

what the monkey is a "number cube"?

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

So there are 6 possible outcomes

jimthompson5910 (jim_thompson5910):

of those 6, how many outcomes are desired outcomes?

OpenStudy (anonymous):

2 (both 5 and 2) OHHHH its 1/3 cause it simplifies!! Ahhh I get it haha

jimthompson5910 (jim_thompson5910):

yep 2/6 = 1/3

OpenStudy (anonymous):

Thank you so much you were a great help! haha I cant believe I forgot that 2/6 = 1/3 !

jimthompson5910 (jim_thompson5910):

you're welcome